# Information conservation in de Sitter tunneling

### v1

@@ Abstract
In this paper, we consider the three most general cases of progressive de Sitter spacetime. The charged and magnetic particles tunnel into the magnetically charged Reissner-Nordström de Sitter black hole (the most general case of a static black hole), the Kerr-Newman-Kasuya de Sitter black hole (the most general case of a rotating black hole), and Bardeen de Sitter black hole (black hole without singularities). We use Parikh-Wilczek method to calculate the radiation spectra of these black holes respectively, and find that they deviate from the pure thermal spectra, satisfying the unitary principle. Our results support the conservation of information and are generally true for all asymptotic de Sitter spacetimes.
@@ 1 Introduction
It is well known that according to the unitary principle, a fundamental principle of quantum mechanics, pure states evolve over time as pure states and cannot become mixed states. But black holes seem to defy this. In 1975, Hawking used quantum field theory in curved spacetime to show that particles can escape black hole through quantum effects 1 . In 1976, Hawking proved that Hawking radiation is a pure thermal spectrum that carries no information 2 . That is, matter that enters the black hole in a pure state radiates out in a mixed state. The information is lost in the process of Hawking radiation, violating the principle of unitarity.
Page argues that Hawking’s calculations did not take into account the effects of quantum entanglement. Although Hawking radiation does not carry information itself, it has been entangled with particles inside the black hole, thus transmitting information from the black hole to the outside. In 1993, Page proposed that in order to satisfy the unitary principle, the entanglement entropy of radiation should satisfy the Page curve 3 . In simple terms, although the radiation is a mixed state, the direct product of the radiation state and the black hole state is still a pure state. In recent years, there have been some breakthroughs in this approach 4 ; 5 ; 6 ; 7 ; 8 ; 9 . But there is an essential difficulty with this approach. Since there is no complete theory of quantum gravity, it is impossible to know the concrete form of the density matrix. The entropy of entanglement is determined by the density matrix:
@@
In 2000, Parikh and Wilczekr believed that Hawking radiation was a tunneling process. They started from the perspective of energy conservation and the change in spatiotemporal background caused by the contraction of black holes after radiation, that is, they considered the self gravity effect during the particle emission process, thus avoiding the difficulty of calculating the specific form of the density matrix. They also came to the conclusion that the radiation process satisfies the unitary principle and information conservation 10 . After the Parikh-Wilczek method was proposed, researchers successively studied various static and stationary rotating black holes and reached the same conclusion. Due to the effect of self gravity, the spatiotemporal background changes, and the radiation spectrum will deviate from the pure thermal spectrum, thus proving that Hawking radiation satisfies the unitary principle. 11 ; 12 ; 13 ; 14 ; 15 ; 16 ; 17 ; 18 ; 19 ; 20 ; 21 ; 22 ; 23 ; 24 ; 25 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 ; 32 ; 33 ; 34 . But until now, apart from the Hawking radiation of the Schwarzschild de Sitter spacetime discussed in ref. 27 , there has been little discussion on the asymptotic de Sitter spacetime that is more similar to the real universe.
In Section 2, we separately calculated the radiation spectra near the event horizon and cosmic horizon of the magnetically charged Reissner-Nordström de Sitter black hole, and briefly discussed our results. In Section 3 and Section 4, we performed similar calculations on the Kerr-Newman-Kasuya de Sitter black hole and Bardeen de Sitter black hole, respectively. Finally, in Section 5, we summarized the results and the discussion of the results obtained in the previous section. We have adopted the system of natural units in this paper ($G\equiv\hbar\equiv c\equiv 1$).
@@ 2 Magnetically charged Reissner-Nordström de Sitter black hole
According to ref. 32 , for the case of both electric and magnetic charges, the electromagnetic field tensor can be defined as:
@@
Where $\alpha$ is a fixed angle. From (3) and (4), we can deduce that:
@@
According to the no-hair theorem, we can consider a black hole as a conductor sphere, and assume that the charge density and magnetic charge density satisfy the following relationship to ensure that (6) and (7) hold true:
@@
The corresponding generalized coordinates for (10) are:
@@
Where $M$ is the mass of the black hole, $Q_{e}$ is the charge of the black hole, $Q_{g}$ is the magnetic charge of the black hole, and $\Lambda$ is the cosmological constant. According to (9), the line element of Magnetically charged Reissner-Nordström de Sitter black hole can be rewritten as:
@@
Clearly, (14) has four solutions: $r_{non}$, $r_{-}$, $r_{+}$, and $r_{c}$. Among these, $r_{non}$ is a negative solution without physical significance, $r_{-}$ represents the inner event horizon of black hole, $r_{+}$ denotes the outer event horizon of black hole, and $r_{c}$ signifies the cosmic horizon.
@@
$\lambda=$ $\left[-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2}+\sqrt{(-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2})^{2}-4(9-36\Lambda Q_{h}^{2})^{3}}\right]^{1/3},$ (19)
@@
$\zeta=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}(1-4\Lambda Q_{h}^{2})}{\Lambda\lambda}-\frac{\lambda}{3(2)^{1/3}\Lambda}-\frac{12M}{\eta}}.$ (21)
@@
$r_{i}=\frac{1}{2}\eta_{i}-\frac{1}{2}\zeta_{i}.$ (22)
Where:
$\lambda_{i}=$ $\left[-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2}+\sqrt{(-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2})^{2}-4(9-36\Lambda Q_{h}^{2})^{3}}\right]^{1/3},$ (23)
$\eta_{i}=$ $\sqrt{\frac{2}{\Lambda}+\frac{3(2)^{1/3}(1-4\Lambda Q_{h}^{2})}{\Lambda\lambda_{i}}+\frac{\lambda_{i}}{3(2)^{1/3}\Lambda}},$ (24)
$\zeta_{i}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}(1-4\Lambda Q_{h}^{2})}{\Lambda\lambda_{i}}-\frac{\lambda_{i}}{3(2)^{1/3}\Lambda}-\frac{12M}{\Lambda\eta_{i}}}.$ (25)
@@
$\lambda_{f}=$$\{-54+972\Lambda(M-\omega)^{2}-648\Lambda(Q_{h}-q_{h})^{2}+$
$\sqrt{[-54+972\Lambda(M-\omega)^{2}-648\Lambda(Q_{h}-q_{h})^{2}]^{2}-4[9-36\Lambda(Q_{h}-q_{h})^{2}]^{3}}\}^{1/3},$ (27)
@@
$\zeta_{f}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}-q_{h})^{2}]}{\Lambda\lambda_{f}}-\frac{\lambda_{f}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega)}{\Lambda\eta_{f}}}.$ (29)
@@
$\lambda^{\prime}=$$\{-54+972\Lambda(M-\omega^{\prime})^{2}-648\Lambda(Q_{h}-q_{h}^{\prime})^{2}+$
$\sqrt{[-54+972\Lambda(M-\omega^{\prime})^{2}-648\Lambda(Q_{h}-q_{h}^{\prime})^{2}]^{2}-4[9-36\Lambda(Q_{h}-q_{h}^{\prime})^{2}]^{3}}\}^{1/3},$ (34)
@@
$\zeta^{\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}-q_{h}^{\prime})^{2}]}{\Lambda\lambda^{\prime}}-\frac{\lambda^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega^{\prime})}{\Lambda\eta^{\prime}}}.$ (36)
@@
$r_{i}^{\prime}=\frac{1}{2}\eta_{i}^{\prime}+\frac{1}{2}\zeta_{i}^{\prime}.$ (37)
Where:
$\lambda_{i}^{\prime}=$ $\left[-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2}+\sqrt{(-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2})^{2}-4(9-36\Lambda Q_{h}^{2})^{3}}\right]^{1/3},$ (38)
$\eta_{i}^{\prime}=$ $\sqrt{\frac{2}{\Lambda}+\frac{3(2)^{1/3}(1-4\Lambda Q_{h}^{2})}{\Lambda\lambda_{i}^{\prime}}+\frac{\lambda_{i}^{\prime}}{3(2)^{1/3}\Lambda}},$ (39)
$\zeta_{i}^{\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}(1-4\Lambda Q_{h}^{2})}{\Lambda\lambda_{i}^{\prime}}-\frac{\lambda_{i}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12M}{\Lambda\eta_{i}^{\prime}}}.$ (40)
@@
$\lambda_{f}^{\prime}=$$\{-54+972\Lambda(M+\omega)^{2}-648\Lambda(Q_{h}+q_{h})^{2}+$
$\sqrt{[-54+972\Lambda(M+\omega)^{2}-648\Lambda(Q_{h}+q_{h})^{2}]^{2}-4[9-36\Lambda(Q_{h}+q_{h})^{2}]^{3}}\}^{1/3},$ (42)
@@
$\zeta_{f}^{\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}+q_{h})^{2}]}{\Lambda\lambda_{f}^{\prime}}-\frac{\lambda_{f}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega)}{\Lambda\eta_{f}^{\prime}}}.$ (44)
@@
$\lambda^{\prime\prime}=$$\{-54+972\Lambda(M+\omega^{\prime})^{2}-648\Lambda(Q_{h}+q_{h}^{\prime})^{2}+$
$\sqrt{[-54+972\Lambda(M+\omega^{\prime})^{2}-648\Lambda(Q_{h}+q_{h}^{\prime})^{2}]^{2}-4[9-36\Lambda(Q_{h}+q_{h}^{\prime})^{2}]^{3}}\}^{1/3},$ (49)
@@
$\zeta^{\prime\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}+q_{h}^{\prime})^{2}]}{\Lambda\lambda^{\prime\prime}}-\frac{\lambda^{\prime\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega^{\prime})}{\Lambda\eta^{\prime\prime}}}.$ (51)

### v2

@@ Abstract
In this paper, we consider the three most general cases of progressive de Sitter spacetime. The charged and magnetic particles tunnel into the magnetically charged Reissner-Nordström de Sitter black hole (the most general case of a static black hole), the Kerr-Newman-Kasuya de Sitter black hole (the most general case of a rotating black hole), and Bardeen de Sitter black hole (black hole without singularities). We use Parikh-Wilczek method to calculate the radiation spectra of these black holes respectively, and find that they deviate from the pure thermal spectra, satisfying the unitary principle. Our results support the conservation of information and are generally true for all asymptotic de Sitter space-times.
@@ 1 Introduction
It is well known that according to the unitary principle, a fundamental principle of quantum mechanics, pure states evolve over time as pure states and cannot become mixed states. But black holes seem to defy this. In 1975, Hawking used quantum field theory in curved spacetime to show that particles can escape black hole through quantum effects [1]. In 1976, Hawking proved that Hawking radiation is a pure thermal spectrum that carries no information [2]. That is, matter that enters the black hole in a pure state radiates out in a mixed state. The information is lost in the process of Hawking radiation, violating the principle of unitarity.
Page argues that Hawking’s calculations did not take into account the effects of quantum entanglement [3]. Although Hawking radiation does not carry information itself, it has been entangled with particles inside the black hole, thus transmitting information from the black hole to the outside. In 1993, Page proposed that in order to satisfy the unitary principle, the entanglement entropy of radiation should satisfy the Page curve [4]. In simple terms, although the radiation is a mixed state, the direct product of the radiation state and the black hole state is still a pure state. In recent years, there have been some breakthroughs in this approach [5, 6, 7, 8, 9, 10]. But there is an essential difficulty with this approach. Since there is no complete theory of quantum gravity, it is impossible to know the concrete form of the density matrix. The entropy of entanglement is determined by the density matrix:
@@
In 2000, Parikh and Wilczekr believed that Hawking radiation was a tunneling process. They started from the perspective of energy conservation and the change in spatiotemporal background caused by the contraction of black holes after radiation, that is, they considered the self gravity effect during the particle emission process, thus avoiding the difficulty of calculating the specific form of the density matrix. They also came to the conclusion that the radiation process satisfies the unitary principle and information conservation [11]. After the Parikh-Wilczek method was proposed, researchers successively studied various static and stationary rotating black holes and reached the same conclusion. Due to the effect of self gravity, the spatiotemporal background changes, and the radiation spectrum will deviate from the pure thermal spectrum, thus proving that Hawking radiation satisfies the unitary principle. [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. But until now, apart from the Hawking radiation of the Schwarzschild de Sitter spacetime discussed in ref. [28], there has been little discussion on the asymptotic de Sitter spacetime that is more similar to the real universe.
The black hole information paradox, first highlighted by Hawking’s seminal work [1, 2], remains one of the most profound challenges at the intersection of quantum mechanics and general relativity. While significant progress has been made in anti-de Sitter (AdS) spacetimes through the AdS/CFT correspondence [36, 37], the de Sitter (dS) universe — characterized by a positive cosmological constant $\Lambda$ — presents a critical yet understudied arena. Observational evidence from Type Ia supernovae [38, 38] and cosmic microwave background measurements [39] strongly supports that our universe is undergoing accelerated expansion, well-modeled by dS geometry. This elevates the study of dS black holes from theoretical curiosity to cosmological necessity. Recent breakthroughs in understanding information recovery, such as the Page curve derivation via semiclassical gravity [8, 9], have predominantly focused on asymptotically AdS black holes. However, fundamental differences between AdS and dS spacetimes — notably the presence of a cosmological horizon and the absence of a globally timelike Killing vector in dS — demand independent verification of unitary principles in realistic cosmological settings. Moreover, existing analyses of dS black holes [28] have largely been restricted to static Schwarzschild-type configurations, neglecting the combined effects of rotation (angular momentum), magnetic charges, and singularity resolution — all of which are theoretically plausible and observationally relevant in astrophysical contexts.
This work systematically addresses these gaps by investigating three general classes of asymptotically dS black holes: 1. Magnetically charged Reissner-Nordström de Sitter black holes — incorporating both electric and magnetic charges to test dyonic effects on information flow; 2. Kerr-Newman-Kasuya de Sitter black holes — extending to rotating configurations where frame-dragging could imprint angular correlations in Hawking radiation; 3. Bardeen de Sitter black holes — probing regular (singularity-free) geometries to determine whether information conservation persists in absence of central singularities. By applying the Parikh-Wilczek tunneling formalism [11] across these diverse scenarios, we demonstrate that the radiation spectrum universally deviates from perfect thermality through entropy-driven suppression $\Gamma\sim e^{\Delta s}$, thereby preserving unitarity. Crucially, our analysis reveals how magnetic charge and rotation introduce novel constraints on information leakage through horizon dynamics — a phenomenon unobservable in static Schwarzschild-AdS systems. These findings not only reinforce the consistency of quantum mechanics in cosmological horizons but also provide critical inputs for the emerging dS/CFT correspondence [40, 41, 42, 43, 44], where boundary unitarity be holographically encoded in bulk entropy transitions. This comprehensive approach significantly advances the field by: Firstly, establishing the robustness of information conservation across geometrically distinct dS black holes; Secondly, bridging the gap between semiclassical tunneling models and holographic principles in realistic cosmologies.
In this paper, We used the Parikh-Wilczek method to calculate tunneling in de Sitter spacetime, taking magnetically charged Reissner-Nordström de Sitter black hole, Kerr-Newman-Kasuya de Sitter black hole, and Bardeen de Sitter black hole as examples. We chose magnetically charged Reissner-Nordström de Sitter black hole because it represents the most general case of a static black hole. Similarly, we chose Kerr-Newman-Kasuya de Sitter black hole because it represents the most general case of a rotating black hole. And we chose Bardeen de Sitter black hole because it represents a black hole without singularities. In Section 2, we separately calculated the radiation spectra near the event horizon and cosmic horizon of the magnetically charged Reissner-Nordström de Sitter black hole, and briefly discussed our results. In Section 3, we performed similar calculation on the Kerr-Newman-Kasuya de Sitter black hole. In Section 4, we calculate the same thing for Bardeen de Sitter black hole. But it is noteworthy that for regular black holes, which have a mass distribution that is not concentrated at a single point and thus do not have a singularity, the first law of black hole thermodynamics will be modified, as described in Ref. [45, 46, 47, 48, 49, 50]. In Section 5, we analyzed the effects of quantum correlations, backreaction corrections, and non-thermal modifications. In Section 6, we compared Parikh-Wilczek method with Hamilton-Jacobi method, entanglement entropy method and GUP-based models. The GUP we mentioned is the Generalized Uncertainty Principle. Finally, in Section 7, we summarized the results and the discussion of the results obtained in the previous section. We have adopted the system of natural units in this paper ($G\equiv\hbar\equiv c\equiv 1$).
@@ 2 Magnetically charged Reissner-Nordström de Sitter black hole
According to Ref. [33], for the case of both electric and magnetic charges, the electromagnetic field tensor can be defined as:
@@
Where $\alpha$ is a fixed angle. From Eq. (3) and Eq. (4), we can deduce that:
@@
According to the no-hair theorem, we can consider a black hole as a conductor sphere, and assume that the charge density and magnetic charge density satisfy the following relationship to ensure that Eq. (6) and Eq. (7) hold true:
@@
The corresponding generalized coordinates for Eq. (10) are:
@@
Where $M$ is the mass of the black hole, $Q_{e}$ is the charge of the black hole, $Q_{g}$ is the magnetic charge of the black hole, and $\Lambda$ is the cosmological constant. According to Eq. (9), the line element of Magnetically charged Reissner-Nordström de Sitter black hole can be rewritten as:
@@
Clearly, Eq. (14) has four solutions: $r_{non}$, $r_{-}$, $r_{+}$, and $r_{c}$. Among these, $r_{non}$ is a negative solution without physical significance, $r_{-}$ represents the inner event horizon of black hole, $r_{+}$ denotes the outer event horizon of black hole, and $r_{c}$ signifies the cosmic horizon.
@@
$\zeta=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}(1-4\Lambda Q_{h}^{2})}{\Lambda\lambda}-\frac{\lambda}{3(2)^{1/3}\Lambda}-\frac{12M}{\eta}},$ (20)
$\lambda=$ $\left[-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2}+\sqrt{(-54+972\Lambda M^{2}-648\Lambda Q_{h}^{2})^{2}-4(9-36\Lambda Q_{h}^{2})^{3}}\right]^{1/3}.$ (21)
@@
$r_{i}=\frac{1}{2}\eta-\frac{1}{2}\zeta.$ (22)
@@
$\zeta_{f}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}-q_{h})^{2}]}{\Lambda\lambda_{f}}-\frac{\lambda_{f}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega)}{\Lambda\eta_{f}}},$ (25)
$\lambda_{f}=$$\left\{-54+972\Lambda(M-\omega)^{2}-648\Lambda(Q_{h}-q_{h})^{2}+\right.$
$\left.\sqrt{[-54+972\Lambda(M-\omega)^{2}-648\Lambda(Q_{h}-q_{h})^{2}]^{2}-4[9-36\Lambda(Q_{h}-q_{h})^{2}]^{3}}\right\}^{1/3}.$ (26)
@@
$\zeta^{\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}-q_{h}^{\prime})^{2}]}{\Lambda\lambda^{\prime}}-\frac{\lambda^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M-\omega^{\prime})}{\Lambda\eta^{\prime}}},$ (32)
$\lambda^{\prime}=$$\left\{-54+972\Lambda(M-\omega^{\prime})^{2}-648\Lambda(Q_{h}-q_{h}^{\prime})^{2}+\right.$
$\left.\sqrt{[-54+972\Lambda(M-\omega^{\prime})^{2}-648\Lambda(Q_{h}-q_{h}^{\prime})^{2}]^{2}-4[9-36\Lambda(Q_{h}-q_{h}^{\prime})^{2}]^{3}}\right\}^{1/3}.$ (33)
@@
$r_{i}^{\prime}=\frac{1}{2}\eta+\frac{1}{2}\zeta.$ (34)
@@
$\zeta_{f}^{\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}+q_{h})^{2}]}{\Lambda\lambda_{f}^{\prime}}-\frac{\lambda_{f}^{\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega)}{\Lambda\eta_{f}^{\prime}}},$ (37)
$\lambda_{f}^{\prime}=$$\left\{-54+972\Lambda(M+\omega)^{2}-648\Lambda(Q_{h}+q_{h})^{2}+\right.$
$\left.\sqrt{[-54+972\Lambda(M+\omega)^{2}-648\Lambda(Q_{h}+q_{h})^{2}]^{2}-4[9-36\Lambda(Q_{h}+q_{h})^{2}]^{3}}\right\}^{1/3}.$ (38)
@@
$\zeta^{\prime\prime}=$ $\sqrt{\frac{4}{\Lambda}-\frac{3(2)^{1/3}[1-4\Lambda(Q_{h}+q_{h}^{\prime})^{2}]}{\Lambda\lambda^{\prime\prime}}-\frac{\lambda^{\prime\prime}}{3(2)^{1/3}\Lambda}-\frac{12(M+\omega^{\prime})}{\Lambda\eta^{\prime\prime}}},$ (44)
$\lambda^{\prime\prime}=$$\left\{-54+972\Lambda(M+\omega^{\prime})^{2}-648\Lambda(Q_{h}+q_{h}^{\prime})^{2}+\right.$
$\left.\sqrt{[-54+972\Lambda(M+\omega^{\prime})^{2}-648\Lambda(Q_{h}+q_{h}^{\prime})^{2}]^{2}-4[9-36\Lambda(Q_{h}+q_{h}^{\prime})^{2}]^{3}}\right\}^{1/3}.$ (45)